OpenStudy (anonymous):

Limits and difference quotient: f(x) = x^2 + x, find f'(3) ...also, find the value of the tangent line for f at x=3

OpenStudy (anonymous):

f'(x) (power ule)

OpenStudy (jhannybean):

Finding the value of the tangent line at x=3 is finding the slope of the function.

OpenStudy (anonymous):

Differentiate each term by term.

OpenStudy (anonymous):

yes.

OpenStudy (jhannybean):

So once you find your derivative, plug in x = 3

OpenStudy (anonymous):

1) find f'(x) 2) find f'(3) by plugging in 3 for x, into the derivative. this will be your slope. then use point slope formula.

OpenStudy (anonymous):

yeah.. find f(2), this is the y of the point.

OpenStudy (jhannybean):

And you'll find the value of the SLOPE of your function at that point. or just the derivative at that point.

OpenStudy (jhannybean):

Is that what he he looking for, the equation of the tangent line?

OpenStudy (jhannybean):

perhaps I misinterpreted the question.

OpenStudy (anonymous):

he said the "value" which is very ambiguous.

OpenStudy (jhannybean):

Yeah, I figured.

OpenStudy (anonymous):

well, at least he knows what he needs, the slope of the tangent line at x=3, or the equation of it... but we gave the assistance accordingly.

OpenStudy (jhannybean):

I'm just waiting for the OP to respond.

OpenStudy (anonymous):

we're studying limits, and getting into derivatives. I think he's looking for us to use the difference quotient.

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

so, I'd calculate the limit x->0 using the equation of the slope at the point given and one we create?

OpenStudy (freckles):

f(x)=x^2+x so f(x+h)=?

OpenStudy (anonymous):

(x+h)^2 + (x + h)

OpenStudy (anonymous):

((x+h)^2 + (x + h) - 12) / h

OpenStudy (jhannybean):

expand (x+h)^2

OpenStudy (anonymous):

x^2 + 2xh + h^2

OpenStudy (jhannybean):

\[f(x) = x^2 + x\]\[f(x+h) = x^2 +2xh + h^2 +x+h\]NOw we need to subtract this by the riginal function.

OpenStudy (jhannybean):

So x^2 +2xh + h^2 +x+h - f(x) = x^2 +2xh +h^2 +x+h - x^2 - x

OpenStudy (jhannybean):

simplify by combining like terms.

OpenStudy (jhannybean):

Then you take the simplified version of this function, and divide by h. By taking the limit as h approaches 0, anything that still has an h in it will converge to 0

OpenStudy (anonymous):

I came up with an answer of 3, does that sound correct?

OpenStudy (jhannybean):

Let's check

OpenStudy (anonymous):

actually, maybe 2x + 3

OpenStudy (anonymous):

2x + 1

OpenStudy (jhannybean):

x^2 +2xh +h^2 +x+h - x^2 - x = 2xh +h^2 +h \[\lim_{x \rightarrow 0} \frac{2xh+h^2+h}{h}\]\[\lim_{x \rightarrow 0} ~(2x+h^2+1)\]\[=2x+1~ \checkmark \]

OpenStudy (jhannybean):

Good, now what do you have to do?

OpenStudy (anonymous):

so, is that the answer to both questions, meaning f prime and the slope of the tangent?

OpenStudy (anonymous):

or is f prime what we came up with, and the slope y=2x + 1

OpenStudy (jhannybean):

oh, this is the equation of your tangent line. \(f'(x) = 2x + 1\). Now we're evaluating it to find the slope at the point x = 3

OpenStudy (anonymous):

oh, and plug in x

OpenStudy (anonymous):

yes, what's the slope there?

OpenStudy (jhannybean):

solve \[f'(3) = 2(3)+1\]

OpenStudy (anonymous):

slope is 2?

OpenStudy (jhannybean):

No, your slope evaluated at x=3 is not 2, unfortunately.

OpenStudy (jhannybean):

Okay. What we found is the equation of the tangent line itself. The slope of that line is 2. Your question also asks to find the value of the slope when x = 3 that means we will have a different tangent line equation at that point. Your question is worded a bit differently than what I am used to, so I was a little confused, sorry.

OpenStudy (anonymous):

so, what is the slope when x=3

OpenStudy (jhannybean):

f'(3) = 2(3)+1 = ?

OpenStudy (jhannybean):

simplify that expression.

OpenStudy (anonymous):

7, so the slope of the tangent line at x=3 is 7

OpenStudy (jhannybean):

You finally got it :)